OpenStudy (anonymous):

If y = 3/5, what is 22y-1 i need answer in deciamal form

OpenStudy (astrophysics):

Plug in 3/5 where y is 22(3/5)-1 = 12.2

OpenStudy (anonymous):

the answer choices are 1.15 .76 .6 1.4

OpenStudy (astrophysics):

I feel as if you're missing information, because that's the right answer.

OpenStudy (anonymous):

no the 2 is sqaured its not 22 i typed in the wrong thing

OpenStudy (anonymous):

if y=3/5 what is 2^2y-1

OpenStudy (astrophysics):

\[2^{2y}-1?\]

OpenStudy (astrophysics):

Like that?

OpenStudy (anonymous):

yea

OpenStudy (astrophysics):

1.29 = 1.3

OpenStudy (astrophysics):

\[2^{2*3/5}-1 = 1.3\]

OpenStudy (anonymous):

thank u

OpenStudy (astrophysics):

Yw